Company X is considering the purchase of a machinery that would result in the following net...

  1. Company X is considering the purchase of a machinery that would result in the following net cash inflows:

    End of Year            1                  2                  3                  4                  5

                                  ----------------------------------------------------------------------------

    Net Cash Inflow       $5,000          4,000            3,000            2,000           1,000   The required rate of return for a project of this risk is 15%. What would the price of the machine have to be for the company just to break-even on the investment?
Answer #1

To break-even, the price of the machine should equal the present value of the cash inflows generated by the machine.

Present value = future value / (1 + required return)number of years

Total present value of cash inflows = $10,985.63

The break-even price of the machine is $10,985.63

A B C Cash 1 Year Inflow 2 1 $5,000 2 $4,000 3 $3,000 4 $2,000 5 $1,000 Total PV of Cash Inflow $4,347.83 $3,024.57 $1,972.55

1 Year 2 1 32 4 3 Cash Inflow 5000 4000 3000 2000 1000 Total PV of Cash Inflow =B2/(1+15%)^A2 =B3/(1+15%)^A3 =B4/(1+15%)^A4 =